# Fast Gradient Computation for RoPE Attention in Almost Linear Time

### v2

@@ 6.5 Fast Computation in Almost Linear Time
Assuming the entries of $A_{1},A_{2},X1,X_{2},Y,E$ are represented using $O(\log n)$ bits, there is an $n^{1+o(1)}$ time algorithm to solve $\mathsf{AAttLGC}(n,d=O(\log n),B=o(\sqrt{\log n}))$, from Def. 3.3, with the accuracy upper bounded by $\frac{1}{\operatorname{poly}(n)}$ To be more specific, a gradient vector $\widetilde{g}\in\mathbb{R}^{d^{4}}$ comes out of our algorithm where $\|\frac{\mathrm{d}\mathsf{Loss}}{\mathrm{d}x}-\widetilde{g}\|_{\infty}\leq\frac{1}{\operatorname{poly}(n)}$.
@@
By Lemma 6.6 and Lemma 6.5, There are matrices $\gamma(x),\gamma_{1}(x)\in\mathbb{R}^{n\times n}$ and $\gamma_{2}(x)$, we have
@@
We assume Lemma 6.5 and Lemma 6.6 are true from Lemma 6.2 to Lemma 6.4. Thus, we can have the following based on Lemma 6.5 Lemma 6.6.
@@
The cost is $n^{1+o(1)}$ time for every Lemma in Lemmas 6.2, 6.3, 6.4, 6.5 and 6.6.
We have the reformulated gradient from Lemma 4.1 as follows.
@@
where the first equality is based on Lemma 4.1, the second equality is due to the distributive law, the third equality derives from the definition of $\ell_{\infty}$ norm, the fourth equality is due to the def. of vectorization, and the fifth inequality derives from the Lemmas in Lemma 6.5 and Lemma 6.6.
@@
7 Hardness
In this section, we introduce our hardness results.
Theorem 7.1 (Lower bound).
Assuming SETH, for any $q>0$, for the $\mathsf{ARAttLGC}(n,d=O(\log n),B=\omega(\sqrt{\log n})$, there does not exist an algorithm which can be executed in time $O(n^{2-q})$ based on Def. 3.3.
@@ 7 Hardness
8 Conclusion
This paper presents the first practical and efficient backward gradient algorithm, assuming bounded entries for the RoPE-based attention mechanism. We achieve almost linear time complexity by leveraging polynomial methods and the Fast Fourier Transform, making the forward and backward computations comparably efficient. Additionally, we demonstrate that conditions exist under which performance better than quadratic can be realized, consistent with the lower bounds suggested by the Strong Exponential Time Hypothesis (SETH). These findings not only improve the computational efficiency of RoPE-based attention mechanisms but also provide a foundation for exploring sub-gradient computations in other advanced attention variants of neural networks. This work highlights the connection between algorithm design and computational complexity theory, unveiling new possibilities for the development of large transformer models. Future research could extend these results to cases involving unbounded entries and assess the real-world implications of these theoretical advancements for large language models. Furthermore, applying this approach to other positional encoding mechanisms could further enhance the scalability of state-of-the-art transformer models.

### v3

@@ D.5 Fast Computation in Almost Linear Time
Theorem D.6 (Main result, Low Rank Approximate RoPE Attention Gradient, Restatement of Theorem 5.7).
Assuming the entries of $A_{1},A_{2},X_{1},X_{2},Y,E$ are represented using $O(\log n)$ bits, there is an $n^{1+o(1)}$ time algorithm to solve $\mathsf{AAttLGC}(n,d=O(\log n),B=o(\sqrt{\log n}))$, from Def. 3.3, with the accuracy upper bounded by $\frac{1}{\operatorname{poly}(n)}$ . To be more specific, a gradient vector $\widetilde{g}\in\mathbb{R}^{d^{4}}$ comes out of our algorithm where $\|\frac{\mathrm{d}\mathsf{Loss}}{\mathrm{d}x}-\widetilde{g}\|_{\infty}\leq\frac{1}{\operatorname{poly}(n)}$.
@@
By Lemma D.5 and Lemma D.4, There are matrices $\gamma(x),\gamma_{1}(x)\in\mathbb{R}^{n\times n}$ and $\gamma_{2}(x)$, we have
@@
We assume Lemma D.4 and Lemma D.5 are true from Lemma D.1 to Lemma D.3. Thus, we can have the following based on Lemma D.4 Lemma D.5.
@@
The cost is $n^{1+o(1)}$ time for every Lemma in Lemmas D.1, D.2, D.3, D.4 and D.5.
We have the reformulated gradient from Lemma C.4 as follows.
@@
where the first equality is based on Lemma C.4, the second equality is due to the distributive law, the third equality derives from the definition of $\ell_{\infty}$ norm, the fourth equality is due to the def. of vectorization, and the fifth inequality derives from the Lemmas in Lemma D.4 and Lemma D.5.
@@
Remark D.7.
The assumption in Theorem 5.7 is practical. In practice, especially in recent long context tasks, the $n$ is large, e.g., $n=2\times 10^{6}$ for Google’s Gemini 1.5 Pro  [13], while the model training uses a half-precision floating-point format, e.g., the bit number is $16$. Furthermore, our assumption is “tight”, where if we slightly weaken the assumption, there is no algorithm that can solve the RoPE attention gradient computation in truly sub-quadratic complexity (Theorem 6.1).
Our Theorem 5.7 accurately approximates ($\epsilon=1/\operatorname{poly}(n)$) the RoPE attention gradient computation in almost linear time $n^{1+o(1)}$ under practical assumptions (see the above Remark D.7). Thus, our methods solve the last puzzle of RoPE attention acceleration. Combined with previous work on RoPE attention inference (see Lemma 5.1), this may make RoPE attention practical as we overcome the theoretical quadratic time complexity barrier both in inference and training.
Appendix E Hardness
In this section, we provide the lower bound results to compute the gradient of RoPE attention.
Theorem E.1 (Lower bound).
Assuming $\mathsf{SETH}$, for any $q>0$, for the $\mathsf{ARAttLGC}(n,d=O(\log n),B=\omega(\sqrt{\log n})$, there does not exist an algorithm which can be executed in time $O(n^{2-q})$ based on Def. 3.3.